The square of a negative number is 35 more than 2 times the negative number. Find the number.
The number =

Answer by ankor@dixie-net.com(22740)

The square of a negative number is 35 more than 2 times the negative number.
Find the number.
let n = the number
n^2 = 2(n) + 35
n^2 = 2n + 35
arrange as a quadratic equation
n^2 - 2n - 35 = 0
Factors to
(n-7)(n+5) = 0
they want the negative solution
n = -5
:
;
See if that works
(-5)^2 = 2(-5) + 35
25 = -10 + 35